OpenStudy (samigupta8):

if a, b and c are three positive real numbers then the minimum value of the expression b+c/a +c+a/b +a+b/c is

OpenStudy (mathstudent55):

Do you mean: \(b + \dfrac{c}{a} + c + \dfrac{a}{b} + a + \dfrac{b}{c} \) (which is what you wrote) or \(\dfrac{b + c}{a} + \dfrac{c + a}{b} + \dfrac{a + b}{c} ~~~~~ ?\)

OpenStudy (sumi29):

@samigupta8 Please use the equation mode to input your question, otherwise it gets kinda hard to understand the question.

OpenStudy (shubhamsrg):

I am assuming it to be the latter one. Simplify your expression like this : (b/a) + (c/a) + (c/b) + (a/b) + (a/c) + (b/c) = ( b/a + a/b) + ( c/a + a/c) + ( c/b + b/c) Doesthis help your case ?

OpenStudy (rsadhvika):

so 3 ?

OpenStudy (rsadhvika):

( b/a + a/b) + ( c/a + a/c) + ( c/b + b/c) >=1 >=1 >=1 >= 3 minimum value = 3

OpenStudy (shubhamsrg):

>=2 right ? so 6 :|

OpenStudy (rsadhvika):

how ? am/gm >= 1 right ?

OpenStudy (shubhamsrg):

(b/a + a/b)/2 >= 1 and hence the conclusion .

OpenStudy (rsadhvika):

yeah missed the /2 !

OpenStudy (samigupta8):

bt how come b/a +a/b>=1

OpenStudy (samigupta8):

kk... i understood